OpenStudy (anonymous):

Calculate the mass of 1 a.m.u in g

sam (.sam.):

1 amu =1/12 of Carbon-12 So, \[1amu=\frac{1}{12} \times [6(1.672622 \times 10^{-24})+ 6(1.674927 \times 10^{-24}+6(9.10938 \times 10^{-28})]\] \[1amu \approx 1.67 \times 10^{-24}grams\]